# HOLMBERG IX: A UNIQUE, INFANT BUT INACTIVE GALAXY as Revealed via a Multiwavelength Approach

### 03 Apr 2026 v1

@@ Appendix A JUSTIFICATION FOR THE ASSUMED STAR FORMATION HISTORY
On the other hand, among all forms of SFHs, the instantaneous SF burst yields the reddest colors for stellar populations at constant age; any more complex SFHs lead to older ages for fixed color indices (detailed analyses of SFH impacts on UV properties for stellar populations are presented in Section 6 in Mao et al., 2012). According to the previously derived age Myr for Holmberg IX (Sabbi et al., 2008; Weisz et al., 2008; Okamoto et al., 2015), we are also able to justify the conclusion about the SFH and the age drawn in this work. If Holmberg IX and its subregions experience more complex SFHs, then the given UV colors require the galaxy in possession of stars older than 250 Myr which should have been detected through the optical-NIR channels.

### 16 Apr 2026 v2

@@ Appendix A JUSTIFICATION FOR THE ASSUMED STAR FORMATION HISTORY
On the other hand, among all forms of SFHs, the instantaneous SF burst yields the reddest colors for stellar populations at a constant age; any more complex SFHs lead to older ages for fixed color indices (detailed analyses of SFH impacts on UV properties for stellar populations are presented in Section 6 in Mao et al., 2012). According to the previously derived age Myr for Holmberg IX (Sabbi et al., 2008; Weisz et al., 2008; Okamoto et al., 2015), we are also able to justify the conclusion about the SFH and the age drawn in this work. If Holmberg IX and its subregions experience more complex SFHs, then the given UV colors require the galaxy in possession of stars older than 250 Myr, which should have been detected through the optical-NIR channels.